In the following problem, check that it is appropriate to use the normal approximation to the binomial. Then use the normal distribution to estimate the requested probabilities.

Do you try to pad an insurance claim to cover your deductible? About 43% of all U.S. adults will try to pad their insurance claims! Suppose that you are the director of an insurance adjustment office. Your office has just received 126 insurance claims to be processed in the next few days. Find the following probabilities. (Round your answers to four decimal places.)

(a) half or more of the claims have been padded
0.0563 (incorrect)

(b) fewer than 45 of the claims have been padded
0.0495 (incorrect)

(c) from 40 to 64 of the claims have been padded
0.9562 (incorrect)

(d) more than 80 of the claims have not been padded

0 (incorrect)

I got all of them wrong and I need help what the answer is.
Answer #1

normal approximation to the binomial can be use only if

np ≥ 5   and n(1 − p) ≥ 5

Here, 126*0.43>5 and 126*0.57>5

Mean = 0.43*126 = 54.18

Std = (126*0.43*0.57)^0.5 = 5.5572

a)

P(x>63)

Step 1: Sketch the curve. The probability that X> 63 is equal to the blue area under the curve. 0.08 y0.04 0.01 50 Step 2: Since μ 54.18 and σ 5.5572 we have. X μ63-54.18 P(X>63 )-P(X-μ>63-5418)-P - Since Z and 63 5418 1,59 we have P(X 63) P1.59) Step 3: Use the standard normal table to conclude that: P (Z > 1.59) 0.0559

b)

P(x<45)

c)

P(40<x<64)

d)

P(x<80) = 1.000